OpenStudy (anonymous):

Solve the following system of equations. 4x + 7y – 2z = 0 3x – 5y + 3z = 9 3x + 6y – z = 1 Choices: A.(1, 0, 2) B.(–1, 0, –2) C. (1, 0, –2) D.(–1, 0, 2) Is the correct answer D?

OpenStudy (anonymous):

Not quite notice when we substitute those values into the 2nd equation it does not equal 9

OpenStudy (anonymous):

For these questions, as multiple choice, it can sometimes be faster just to check the answers in the equations and see which one provides true statements every time.

OpenStudy (anonymous):

so is it B?

myininaya (myininaya):

no

myininaya (myininaya):

just do what simple said see which choice holds for the system

myininaya (myininaya):

plug in

OpenStudy (anonymous):

ok

myininaya (myininaya):

what happens when you check A?

myininaya (myininaya):

do each of the equations hold true for (x,y,z)?

myininaya (myininaya):

(x=1,y=0,z=2)*?

OpenStudy (anonymous):

all i did was subract the second and and third equation

OpenStudy (anonymous):

and i got -11y+4z=8

myininaya (myininaya):

ok it is easier if you just plug in those values and see if each of the equations hold

OpenStudy (anonymous):

oh so the naswer is A right?

OpenStudy (anonymous):

answer*

myininaya (myininaya):

right becasuse 4(1)+7(0)-2(2)=4-4=0 both sides = 3(1)-5(0)+3(2)=3+6=9 both sides = 3(1)+6(0)-2=3-2=1 both sides =

OpenStudy (anonymous):

ooooh i get it thank you so much for your help i really appreciate it

OpenStudy (anonymous):

good work. nice finish too myininaya

OpenStudy (anonymous):

thank both so much